OpenStudy (anonymous):

what is the y intercept of the line 5x-2y=20

OpenStudy (igreen):

Put it in the form of y = mx + b. 5x - 2y = 20 -2y = 5x + 20 y = -5/2x - 10

OpenStudy (texaschic101):

iGreen is correct....put it in y = mx + b form, and the number in the b position is the y intercept

OpenStudy (texaschic101):

wait....there's a mistake. @iGreen ....check your math again...although the mistake is not gonna change the answer for the y intercept

OpenStudy (igreen):

My slope?

OpenStudy (texaschic101):

yep.....subtract 5x from both sides

OpenStudy (igreen):

Yeah, it's positive..not negative..

OpenStudy (igreen):

Positive 5/2, not negative..

OpenStudy (texaschic101):

it does not change the answer...the y intercept is still the same

OpenStudy (igreen):

I know..